# IMPACT[Uncaptioned image]: Intelligent Motion Planning with Acceptable Contact Trajectories via Vision-Language Models

### v1

@@ V-A Experiment Setup in Simulation
We build and test our pipeline using PyBullet simulation [41]. In simulation, we create 20 scenes designed to test contact-rich manipulation. These use a hybrid object dataset, which includes tall and bulky items (e.g., sugar boxes and water pitchers) from the YCB [42] data and fragile objects (e.g., wine glasses and stacked bowls) using 3D models generated from TRELLIS [43]. TRELLIS allows constructing delicate object meshes from single-view images and introduces fragility constraints not present in YCB objects. All objects are on a shelf so that the robot cannot reach the target from above. See Fig. 3 for an example scene.
We use three cameras which capture the scene for cost map generation: one in front of the shelf and two on the sides. The camera in front of the shelf also captures the scene image $I$ used as input to GPT. To prevent interactions with the shelf or tabletop, we explicitly assign their costs to be 10. We set the maximum allowed path cost to 10 to balance efficiency in path planning and contact avoidance. To quantify the benefit of using GPT-generated costs, we compare with two control groups: (i) with the maximum path cost set to 0 to obtain a collision-free path, which we test in our “Collision-Free Planning” baseline (see Sec. V-C), and (ii) with the cost of all objects set to 0 to see how the planner will plan the path if all collisions are allowed (see Sec. V-F).
@@ V-B Motion Planning Methods
As reviewed in Sec. IV-B, our method is IMPACT+RRT*. In simulation, we also test IMPACT with RRT [7] and MPC [44] as the motion planners, and thus call the respective methods IMPACT+RRT and IMPACT+MPC. We slightly modify these planning algorithms to improve their effectiveness for manipulation in dense clutter. For example, to encourage the robot to avoid high-cost obstacles, we define a threshold for object cost. During planning, we check the distance to each obstacle exceeding this cost threshold to make sure the nodes close to them are not considered as waypoints of the trajectory. We apply this modification to both RRT and RRT*. For MPC, the cost of a state is defined as the sum of (i) the distance to the target object, (ii) the cost of the current end effector position decided by the cost map and (iii) a penalty for collisions with high-cost obstacles.
@@ V-C1 Collision-Free Planning
This avoids all collisions, and serves as a baseline to demonstrate that in cluttered environments, a collision-free path may not exist. We test this with MPC, RRT, and RRT*.
@@ V-D Evaluation Metrics
Category Path Planning Reach Path Contact High Cost Object Success
Collision Free MPC 76.93% - 12.5 8.05 20.75%
RRT 78.92% - 9.39 6.36 34.75%
RRT* 83.12% - 7.57 6.32 37.75%
IMPACT MPC 81.45% 11.59 12.0 7.97 41.00%
RRT 84.67% 8.52 7.64 6.11 58.75%
RRT* (Ours) 82.50% 8.85 7.50 6.08 63.25%
- LAPP 41.18% - 9.53 10.58 35.00%
TABLE I: Comparison of path planning algorithms and results in PyBullet simulation. We report 5 quantitative metrics (see Sec. V-D). “Reach Target" reports when the robot’s end effector reaches the target after executing the trajectory, and where collisions are allowed. “Success Rate" only counts the trajectories that strictly contain no collisions to (human-designated) high-cost obstacles. The arrow $\uparrow$ indicates larger values of the metric correspond to better performance, and $\downarrow$ represents the opposite. Collision-Free baselines and LAPP do not use GPT to generate object costs, so they do not have values for “Path Cost" in this table.
To evaluate the quality of trajectories in our densely cluttered environments, we compute the following metrics. The robot’s trajectory terminates when the robot’s end-effector reaches the target, or if a time limit is reached.
  • reach_target: whether the robot reaches the target object at any point during the trajectory:
[…]
  • path_cost: the sum of the cost of all collided obstacles during execution:
[…]
  • displacement: displacement of each object:
Here, $e$ is the robot end effector. For each object $o\in\mathcal{O}$, $\mathbf{p}_{o}$ and $\mathbf{p}^{\prime}_{o}$ denote its initial and final position, respectively; the final position considers early termination. The cost of object $o$ is represented by $c_{o}$ while $t_{o}$ is the contact duration between the robot and $o$. After calculating these metrics, we define a trajectory as a success if the following are all true: (i) the robot reaches the target with path_cost $<10$, (ii) contact_duration $<100$ and (iii) for all high-cost obstacles, we have displacement $<0.1$. To decide on which objects are “high-cost” in a given scene, a skilled human annotator pre-selects the 1-2 highest-cost objects in a scene, and those are set for all methods evaluated.
@@ V-D1 User Study Evaluation
While the prior quantitative metrics evaluate trajectory cost, they may not fully capture whether a trajectory is “acceptable” to humans. For example, if all collisions with objects are counted and penalized equally, a trajectory where the robot solidly collides with only one object has a lower cost compared to one where it gently contacts multiple objects. Thus, motion planning algorithms optimize to select the former trajectory, even though the latter trajectory may be more acceptable to humans due to the gentle contacts. Furthermore, tolerance for contact may vary from person to person.
To better assess IMPACT’s ability to generate semantically acceptable behavior, we conduct a user study. We develop a website to collect human feedback, where participants evaluate robot trajectories by watching videos. Each question presents two videos of robot trajectories in the same scene with the same target object. One video is generated from our algorithm, and the other by a baseline algorithm. Users are not informed which algorithm produced each trajectory. Following a similar approach as Mirjalili et al. [46], participants are asked to select the trajectory they find most acceptable. See Fig. 4 for a visualization of the website interface. Our study has been approved by the Institutional Review Board (IRB) at the University of Southern California. None of the human evaluators is an author of this paper.
@@ V-E Simulation Results
Figure 5: Examples of trajectories planned using different motion planning algorithms and cost configurations in PyBullet simulation [41]; our method is IMPACT+RRT* (top row). The scene contains three obstacles: a spray bottle, a power drill and a wine glass. The target object is the jar behind the obstacles. The paths planned by different methods are shown in an overlaid green curve in each image. We also provide LAPP with a language instruction “Can collide with the spray bottle and the power drill." See Sec. V-E for more details.
Table I reports our simulation results. IMPACT+RRT* achieves the highest success rate with lower path cost, shorter contact duration, and smaller displacement of high-cost objects. IMPACT significantly improves the success rate of different motion planning algorithms compared to collision-free planning, while reach_target is similar. This indicates that IMPACT can guide motion planning algorithms to plan trajectories with more acceptable contacts. Furthermore, RRT* achieves the best performance compared to other planning methods. The other baseline, LAPP, has a lower reach_target rate and success rate, mainly because it fails to find a path in some scenes.
Fig. 5 shows trajectories planned by different methods in the same scene. Our method IMPACT+RRT* finds a trajectory that makes contact with the spray bottle. However, the path planned by Collision-free+RRT* intends to avoid all the obstacles, but the robot collides with the fragile wine glass during execution. This happens because the Collision-Free baseline prioritizes moving through the gap between the obstacles and brings the robot too close to the wine glass, increasing the risk of collision. LAPP also guides the robot to make contact with the spray bottle while avoiding the wine glass, but the spray bottle is pushed along the path, preventing the robot from reaching the target. As shown in Table I, this common behavior leads to a higher average high-cost object displacement.
Figure 6: Human evaluation results of simulation experiments. Each bar represents the results of comparing IMPACT+RRT* (i.e., our method) versus one baseline. It is divided into three segments: IMPACT+RRT*, the baseline method, and “Cannot decide." Each segment counts the number of questions where people prefer the trajectory generated by the corresponding method. For example, the top row reports eight questions that compare the trajectory from LAPP and IMPACT+RRT*. Among six of these questions, more participants prefer the trajectory planned by IMPACT+RRT*. In one question, they prefer the trajectory generated by LAPP. In another question, most people do not have a strong preference.
Fig. 6 shows the human evaluation results of simulation experiments. We collect feedback from 27 participants, each of them answering 43 questions. Each question compares a video pair that contrasts our method with some baseline. The results show that IMPACT+RRT* is the most preferred method across all scenes, suggesting that our method produces motion plans that better align with human preferences by leveraging commonsense knowledge in GPT.
@@ V-F Ablation Study
To investigate the benefit of IMPACT, we conduct an ablation study by setting all object costs to 0, so all collisions are allowed during path planning. Results in Table II demonstrate that IMPACT improves the trajectory success rate (see Sec. V-D) of reaching tasks. For all three motion planning algorithms, IMPACT takes properties of obstacles into account and assigns object costs based on different scenes. This encourages motion planning algorithms to avoid obstacles and generate trajectories with reduced contact.
Path Planning Algorithm Cost Success Rate
RRT* IMPACT (Ours) 63.25%
Same Cost for All 42.75%
RRT IMPACT (Ours) 58.75%
Same Cost for All 45.50%
MPC IMPACT (Ours) 41.00%
Same Cost for All 39.50%
TABLE II: Results of our ablation study on object costs. “Same cost for all" refers to assigning all object costs the same value 0 instead of querying GPT to generate costs. See Sec. V-F for details.
@@ VI-A Experiment Setup
Figure 7: Examples of successful trajectories with acceptable contact in the real world. The targets in each scene (from top to bottom) are the tomato can, the tomato, and the matcha can, respectively.
[…]
We mount two Intel RealSense L515 cameras on both sides of the scene to capture RGBD images and generate voxel grids. In the real world, we need to get the position of each object to assign the associated object cost to those voxels. Unlike in simulation, we use Grounded SAM 2 [47, 38, 48, 48, 49, 50] to generate the segmented point cloud of each object as we build the 3D cost map. Each image, along with the object list, is provided as input to Grounded SAM 2, to predict the segmentation mask of the objects. We then generate a multi-view segmented point cloud and convert it to the 3D cost map. The motion planning part is the same as in simulation.
For physical experiments, we test IMPACT+RRT* and LAPP. To adapt LAPP for real-world experiments, we fine-tune it on a small dataset containing 20 real-world RGB images, manually annotated with collision scores. The images are captured by an Intel RealSense D435 camera that faces both the robot and the scene. The annotations explicitly label which objects are safe to collide with (for example, “can collide with plastic bottle”). Objects not mentioned in the language prompts are treated as unsafe by default, aligning with LAPP’s methodology of explicit collision rules. Human operators pair each scene with joint configurations and task-aligned prompts, reflecting the original work’s use of free-form language constraints [10]. This retains LAPP’s pre-trained reasoning while grounding predictions in real-world spatial relationships and safety priorities. Of the 10 test scenes, 9 use objects seen during fine-tuning, while 1 scene contains novel objects to evaluate generalization.
@@
IMPACT+RRT* (Ours) 63%
LAPP 54% (Seen Obj.) 50% (Unseen Obj.)
TABLE III: Results of our method versus LAPP in the real world. Seen objects refers to objects seen during fine-tuning of LAPP while unseen objects are novel objects to LAPP. Our method is zero-shot, so all the objects can be considered unseen to it.
@@ VI-B Real World Results
Table III shows real world results of different methods. IMPACT+RRT* outperforms LAPP in all scenes. Furthermore, the performance gap is larger in the scene with unseen objects (70% versus 50% success). The results suggest the strong generalization ability of our method since it does not need sim2real transfer. However, LAPP requires fine-tuning on unseen real world objects to achieve comparable results.
Failure Cases. While IMPACT generates semantically-acceptable trajectories in most trials, failures can happen during execution. Fig. 8 shows two examples of failed trajectories. The main failure cases are as follows: (i) part of the robot gets stuck on an obstacle, (ii) the target is displaced when the robot pushes an obstacle, and (iii) GPT predicts object costs that are misaligned with human preferences.
Figure 8: Examples of real-world failures. In the left image, the robot gets stuck on the box while reaching the target toy dinosaur. In the right image, the robot collides with the glass bottle. The glass bottle hits the target tomato as it falls and the tomato rolls away.
@@ VII Limitations
While promising, IMPACT has several limitations that point to interesting directions for future work. First, after selecting the most semantically-acceptable trajectory, the robot follows it open-loop. This means it cannot react to unexpected disturbances in real-time. Second, IMPACT relies on having relatively complete RGBD observations to provide to the VLMs. Thus, it may be less effective under partial observability with severe occlusions. Developing a closed-loop procedure that can actively perceive the environment may address these limitations.
@@ VIII Conclusion
In this work, we introduce IMPACT, a framework for motion planning in cluttered environments that leverages the broad knowledge in vision-language models (VLMs) to assess object contact tolerance. IMPACT represents this information in a 3D cost map for motion planning. Our results in simulation and in real-world experiments demonstrate that robots can efficiently reach targets while making semantically-acceptable contact when needed. We hope that this work inspires future work towards flexible and contact-rich robot manipulation in densely cluttered environments.

### v2

@@ V-A Experiment Setup in Simulation
We build and test our pipeline using PyBullet simulation [6]. In simulation, we create 20 scenes designed to test contact-rich manipulation. These use a hybrid object dataset, which includes tall and bulky items (e.g., sugar boxes and water pitchers) from the YCB [3] dataset and fragile objects (e.g., wine glasses and stacked bowls) using 3D models generated from TRELLIS [40]. TRELLIS allows constructing delicate object meshes from single-view images and introduces fragility constraints not present in YCB objects. All objects are placed on a tabletop.
We use three cameras to capture the scene for cost map generation: one in front of the scene and two on the sides. The camera in front also captures the scene image $I$ used as input to the VLM. To prevent interactions with the tabletop, we explicitly assign its cost to be 10. To quantify the benefit of using VLM-generated costs, we compare with two control groups: (i) with no push action allowed to obtain a collision-free path, which we test in our “Collision-Free Planning” baseline (see Sec. V-C), and (ii) with the cost of all objects set to 0 to see how the planner will plan the path if all collisions are allowed (see Sec. V-F).
Category Path Planning Reach Path Contact Unsafe Object Success
Collision Free RRT 66.75% - 1.12 6.47 28.75%
RRT* 61.00% - 1.25 7.78 26.50%
A* w/o direction 20.00% - 8.33 6.51 15.00%
A* 23.33% - 5.14 1.93 15.75%
VLM Cost RRT 62.91% 9.04 1.22 7.84 57.25%
RRT* 59.50% 9.11 1.26 7.80 50.75%
A* w/o direction 58.25% 6.71 5.93 3.82 57.50%
A* (IMPACT) 78.00% 5.93 4.18 2.51 73.75%
- LAPP 50.00% - 9.37 9.22 25.00%
TABLE I: Comparison of path planning algorithms and results in PyBullet simulation. We report five quantitative metrics (see Sec. V-D). “Success Rate” only counts the trajectories that reach the target and satisfy thresholds on “Path Cost,” “Contact Duration,” and “Unsafe Object (human-designated) Displacement.” The arrow $\uparrow$ indicates larger values of the metric correspond to better performance, and $\downarrow$ represents the opposite. Collision-free baselines and LAPP [41] do not use VLMs to generate object costs, so they do not have values for “Path Cost” in this table.
@@ V-B Motion Planning Methods
As presented in Sec. IV-B, IMPACT performs directional contact-aware path planning on the flattened cost map $M^{\prime}$. In simulation, we also test the 3D cost map $V$ with RRT [20] and RRT* [14] as the motion planners, and thus call the respective methods VLM Cost+RRT and VLM Cost+RRT*. Each trajectory consists of a sequence of robot end-effector positions. The total cost of a trajectory $\tau$ is the sum of costs of obstacles the robot arm collides with, $Cost(\tau)=\sum_{i=1}^{|\tau|}V[\tau_{i}]$}. Here, $|\tau|$ denotes the length of the trajectory $\tau$, and $\tau_{i}$ is the position of the $i$-th waypoint. In addition, we test another planning method that directly applies A* search without any directional push safety analysis. We refer to it as VLM Cost+A* w/o direction. It only relies on the isotropic costs in the flattened cost map $M$.
@@ V-C1 Collision-Free Planning
This avoids all collisions, and serves as a baseline to demonstrate that in cluttered environments, a collision-free path may not exist. We test this with A*, A* w/o direction, RRT, and RRT*.
@@ V-D Evaluation Metrics
To evaluate the quality of trajectories in our densely cluttered environments, we compute the following metrics in 3D space. The robot’s trajectory terminates when the robot’s end-effector reaches the target, or if a time limit is reached.
  • reach_target: whether the robot reaches the target object at the end of the trajectory:
    $\texttt{reach\_target}=\|\mathbf{p}^{\prime}_{e}-\mathbf{p}^{\prime}_{o_{\rm targ}}\|<0.01\text{m}.$
  • path_cost: the sum of VLM-assigned costs of all collided obstacles during execution:
    $\texttt{path\_cost}=\sum\nolimits_{o\in\mathcal{O}\setminus o_{\text{targ}}}c_{o}.$
    $\texttt{contact\_duration}=\sum\nolimits_{o\in\mathcal{O}\setminus o_{\text{targ}}}t_{o}.$
  • displacement: the displacement of each object:
Here, $e$ is the robot end effector and $\mathbf{p}^{\prime}_{e}$ is its final 3D position. For each object $o\in\mathcal{O}$, $\mathbf{p}_{o}$ and $\mathbf{p}^{\prime}_{o}$ denote its initial and final 3D position, respectively. The cost of object $o$ is represented by $c_{o}$ while $t_{o}$ is the contact duration between the robot and $o$. After calculating these metrics, we define a trajectory as a success if the following are all true: (i) the robot reaches the target with path_cost <10<10, (ii) contact_duration $100\text{s}$ and (iii) for all unsafe obstacles, we have displacement $10\text{cm}$. To decide on which objects are unsafe in a given scene, a skilled human annotator pre-selects the 1-2 highest-cost objects in a scene, and those are set for all methods evaluated.
@@
Figure 5: Examples of trajectories planned by IMPACT (top row) and LAPP (bottom row) in PyBullet simulation [6]. The obstacles are: a coke bottle, a pitcher, a sugar box and a pile of bowls. The target object is the mug behind the obstacles. The planned paths are shown in an overlaid green curve in each image. We also provide LAPP with a language instruction “Can collide with the pitcher and the sugar box.” See Sec. V-E for more details.
@@ V-D1 User Study Evaluation
While the prior quantitative metrics evaluate trajectory cost, they may not fully capture whether a trajectory is “acceptable” to humans. For example, if all collisions with objects are counted and penalized equally, a trajectory where the robot solidly collides with only one object can have a lower cost compared to one where it gently contacts multiple objects. Thus, motion planning algorithms optimize to select the former trajectory, even though the latter trajectory may be more acceptable to humans due to the gentle contacts. Furthermore, tolerance for contact may vary from person to person.
To better assess IMPACT’s ability to generate semantically acceptable behavior, we conduct a user study approved by the IRB office of our institution. We develop a website to collect human feedback, where participants evaluate robot trajectories by watching videos (see Fig. 4). Each question presents two videos of robot trajectories in the same scene with the same target object. One video is generated by an algorithm that utilizes VLM-assigned costs, and the other by the corresponding Collision-Free baseline with the same planner or by LAPP. Users are not informed which algorithm produced each trajectory. Following a similar approach as Mirjalili et al. [27], participants are asked to select the trajectory they find most acceptable.
@@ V-E Simulation Results
Table I reports our simulation results. IMPACT achieves the highest success rate with lower path cost, shorter contact duration, and smaller displacement of unsafe objects (pre-selected by human). Paths planned by Collision-Free baselines can also lead to unexpected collisions. These baselines prioritize avoiding all obstacles, but the robot may lack continuous feasible joint configurations to execute the path and becomes too close to obstacles, increasing the risk of collisions. IMPACT also outperforms the VLM Cost+A* w/o direction, which we attribute to the contact-aware information embedded in the anisotropic cost map. The LAPP baseline has a lower reach_target and success rate, mainly because it sometimes fails to find paths or push obstacles along the path.
Figure 6: Human evaluation results. The first five rows report results from simulation, and the last row show the real-world experiments. Each bar represents the results of comparing two methods. It is divided into three segments: the left method, the right method, and “Cannot decide.” Each segment counts the number of questions where people prefer the trajectory generated by the corresponding method. For example, in the top row (VLM Cost+RRT vs. Collision-Free+RRT), more participants prefer the trajectory planned by VLM Cost+RRT in 11 cases out of 20, Collision-Free+RRT in six cases, and have no preference in three cases.
Fig. 5 shows trajectories planned by different methods in the same scene. IMPACT finds a trajectory that makes contact with the pitcher. LAPP also guides the robot to make contact with the pitcher while avoiding the pile of bowls and the coke bottle, but the pitcher is pushed along the path, preventing the robot from reaching the target. As shown in Table I, this common behavior leads to a higher average unsafe object displacement.
Fig. 6 shows the human evaluation results collected from 25 participants in total. We collect 100 video pairs in simulation and uniformly sample 20 questions for each participant. Each question contrasts a trajectory generated by one method against some baseline. The results show that the method using VLM-assigned costs is the most preferred method across all scenes, suggesting that cost maps produce motion plans that better align with human preferences by leveraging commonsense knowledge in the VLM.
@@ V-F Ablation Study
To investigate the benefit of VLM-generated object costs, we conduct an ablation study by setting all object costs to 0, so all collisions are allowed during path planning. Results in Table II demonstrate that VLM-assigned costs improve the trajectory success rate (see Sec. V-D) of reaching tasks. For all three motion planning algorithms, the VLM takes properties of obstacles into account and assigns object costs based on different scenes. This encourages planners to avoid obstacles and generate trajectories with reduced contact.
Path Planning Cost Success Rate
Algorithm
RRT VLM Cost 57.25%
Same Cost for All 44.50%
RRT* VLM Cost 50.75%
Same Cost for All 42.50%
A* VLM Cost (IMPACT) 73.75%
Same Cost for All 65.00%
TABLE II: Results of our ablation study on object costs. “Same cost for all” refers to assigning all object costs to be 0, instead of querying the VLM to generate costs. See Sec. V-F for details.
Figure 7: Examples of successful trajectories from IMPACT in the real world. Top row: the robot reaches the target pickleball while moving closer to the soft toy hippo. Bottom row, the robot rotates its gripper to avoid contact with tall bottles and reaches the target matcha can.
@@ VI-A Experiment Setup
IMPACT 61%
LAPP 49% (Seen Obj.) 40% (Unseen Obj.)
TABLE III: Results of our method versus LAPP in the real world. Seen objects refer to objects seen during fine-tuning of LAPP while unseen objects are novel objects to LAPP. Our method is zero-shot, so all the objects can be considered unseen to it.
We evaluate IMPACT on a real robotic system to validate our findings from simulation. This does not involve sim-to-real transfer, as we provide real-world images directly to the VLM for IMPACT. The hardware setup consists of a Franka Panda arm with a standard parallel-jaw gripper, and a flat tabletop surface. A human operator arranges multiple objects in close proximity on top of the surface. For reasonably fair comparisons among methods, the human tries to place objects in consistent locations for each scene.
We mount two Intel RealSense D435 cameras on both sides of the scene to capture RGBD images and generate voxel grids. In the real world, we need to get each object’s position to assign the associated object cost to those voxels. Unlike in simulation, we use Grounded SAM 2 [31, 32] to generate the segmented point cloud of each object as we build the cost map. Each image and the object list are provided as input to Grounded SAM 2 to predict the segmentation masks of the objects. We then get a multi-view segmented point cloud and convert it to the cost map. The motion planning part is the same as in simulation.
For physical experiments, we test IMPACT and LAPP. To adapt LAPP for real-world experiments, we fine-tune it on a small dataset containing 25 real-world RGB images, manually annotated with collision scores. The images are captured by an Intel RealSense D435 camera that faces both the robot and the scene. The annotations explicitly label which objects are safe to collide with (for example, “can collide with plastic bottle”). Objects not mentioned in the language prompts are treated as unsafe by default, aligning with LAPP’s methodology of explicit collision rules. Human operators pair each scene with joint configurations and task-aligned prompts, reflecting the original work’s use of free-form language constraints [41]. This retains LAPP’s pre-trained reasoning while grounding predictions in real-world spatial relationships and safety priorities. Of the 10 test scenes, 8 use objects seen during fine-tuning, while 2 scene contains novel objects to evaluate generalization.
@@ VI-B Real World Results
Table III shows real world results of different methods. IMPACT outperforms LAPP on average across all scenes. Furthermore, the performance gap is larger in scenes with unseen objects (61% versus 40% success). The results suggest the strong generalization ability of our method since it does not need sim-to-real transfer. However, LAPP requires fine-tuning on unseen real world objects to achieve comparable results. We also conduct a user study on real-world experiments. The same 25 participants evaluate 10 pairs of real-world trajectories. As shown in the last row of Fig. 6, IMPACT is preferred in more scenarios. Fig. 7 shows two examples of successful trajectories planned by IMPACT.
Failure Cases. While IMPACT generates semantically-acceptable trajectories in most trials, failures can happen during execution. Fig. 8 shows two examples of failed trajectories. The main failure cases are: (i) part of the robot gets stuck on an obstacle, (ii) the gripper collides with an obstacle before or during rotation, and (iii) the VLM predicts object costs that are misaligned with human preferences.
Figure 8: Examples of real-world failures generated by IMPACT. In the left image, the robot gets stuck on the toy bear while reaching the target glue bottle. In the right image, the robot does not rotate the gripper in time and collides with the ranch bottle.
@@ VII Limitations
While promising, IMPACT has several limitations that point to interesting directions for future work. First, after selecting the most semantically-acceptable trajectory, the robot follows it open-loop. This means it cannot react to unexpected disturbances in real-time. Second, IMPACT relies on having relatively complete RGBD observations, and may be less effective under partial observability with severe occlusions. Developing a closed-loop procedure that can actively perceive the environment may address these limitations.
@@ VIII Conclusion
In this work, we introduce IMPACT, a framework for motion planning in clutter that leverages the broad knowledge in Vision-Language Models (VLMs) to assess object contact tolerance. IMPACT represents this information in an anisotropic cost map for motion planning. Our results in simulation and in real-world experiments demonstrate that robots can efficiently reach targets while making semantically-acceptable contact when needed. We hope that this work inspires future work towards flexible and contact-rich robot manipulation in densely cluttered environments.